# On $\mathcal{NP}$ versus ${\rm co}\mathcal{NP}$ and Frege Systems

### v7

@@ Abstract
We prove in this paper that there is a language $L_{d}$ accepted by some nondeterministic Turing machines but not by any ${\rm co}\mathcal{NP}$-machines (defined later). Then we further show that $L_{d}$ is in $\mathcal{NP}$, thus proving that $\mathcal{NP}\neq{\rm co}\mathcal{NP}$. The main techniques used in this paper are lazy-diagonalization and the novel new technique developed in author’s recent work [Lin21]. Further, since there exists some oracle $A$ such that $\mathcal{NP}^{A}={\rm co}\mathcal{NP}^{A}$, we then explore what mysterious behind it and showing that if $\mathcal{NP}^{A}={\rm co}\mathcal{NP}^{A}$ and under some rational assumptions, the set of all polynomial-time co-nondeterministic oracle Turing machines with oracle $A$ is not enumerable, thus showing that the technique of lazy-diagonalization is not applicable for the first half of the whole step to separate $\mathcal{NP}^{A}$ from ${\rm co}\mathcal{NP}^{A}$. As a by-product, we reach the important result that $\mathcal{P}\neq\mathcal{NP}$ [Lin21] once again, which is clear from the above outcome and the well-known fact that $\mathcal{P}={\rm co}\mathcal{P}$. Next, we show that the complexity class ${\rm co}\mathcal{NP}$ has intermediate languages, i.e., there exists language $L_{inter}\in{\rm co}\mathcal{NP}$ which is not in $\mathcal{P}$ and not ${\rm co}\mathcal{NP}$-complete. We also summarize other direct consequences implied by our main outcome such as $\mathcal{NEXP}\neq{\rm co}\mathcal{NEXP}$ and other which belongs to the area of proof complexity. Lastly, we show a lower bounds result for Frege proof systems, i.e., no Frege proof systems can be polynomial bounded.
@@ 8.1 Lower Bounds for Frege Systems
The following theorem gives a necessary and sufficient condition for the complexity class $\mathcal{NP}$ being closed under complementation. For completeness, the proof is also quoted:
@@ 9 Conclusions
thus resolving a very important and long-standing conjecture in computational complexity theory. Specifically, we first showed that the set of all ${\rm co}\mathcal{NP}$-machines is enumerable. Then, we constructed a universal nondeterministic Turing machine $D$ to simulate all ${\rm co}\mathcal{NP}$-machines and to do the lazy-diagonalization operation to produce a language $L_{d}$ which is not in ${\rm co}\mathcal{NP}$. Next, we applied the technique developed in author’s recent work [Lin21] to show that the language $L_{d}$ accepted by the universal nondeterministic Turing machine $D$ is in fact in $\mathcal{NP}$, thus establishing the desired main conclusion of this paper.
We stress that one of the important techniques used to prove lower bounds (i.e., Theorem 4.1) is the so-called lazy-diagonalization [FS07, AB09], and the essential technique used to show the upper bounds result, i.e., the result that the language $L_{d}\in\mathcal{NP}$, is essentially the same as the one developed in the author’s recent work [Lin21].
We have broken the so-called Relativization Barrier. Since by the result [BGS75], there exists some oracle $A$ such that
@@
which suggests that ordinary diagonalization techniques (lazy-diagonalization) are not capable of proving $\mathcal{NP}\neq{\rm co}\mathcal{NP}$, see e.g. [BGS75]. We explored the mysterious behind this by proving that if $\mathcal{NP}^{A}={\rm co}\mathcal{NP}^{A}$ and under some rational assumptions, then the set of all ${\rm co}\mathcal{NP}$-machines with oracle $A$ is in fact not enumerable, thus showing that the ordinary diagonalization techniques (i.e., lazy-diagonalization) will generally not apply to the relativized versions of the $\mathcal{NP}$ versus ${\rm co}\mathcal{NP}$ problem.
As a by-product, the important result that $\mathcal{P}\neq\mathcal{NP}$ which is shown in [Lin21], also follows. Furthermore, the conclusion of $\mathcal{NEXP}\neq{\rm co}\mathcal{NEXP}$ follows from the similar proof of that $\mathcal{NP}\neq{\rm co}\mathcal{NP}$.
Note again that by Corollary 3 we in fact have shown $\mathcal{P}\neq{\rm co}\mathcal{NP}$ because if $\mathcal{P}={\rm co}\mathcal{NP}$ then $\mathcal{P}={\rm co}\mathcal{P}=\mathcal{NP}$, a contradiction to Corollary 3. As a matter of fact, we have also shown by Theorem 5 that there exists ${\rm co}\mathcal{NP}$-intermediate language in ${\rm co}\mathcal{NP}$ which is not in $\mathcal{P}$ nor ${\rm co}\mathcal{NP}$-complete, i.e., the complexity class ${\rm co}\mathcal{NP}$ also has a rich structure. We have also obtained other interesting consequences such as Corollary 6 which states that there is no super proof system.
As mentioned in the introduction Section, the problem of $\mathcal{NP}$ versus ${\rm co}\mathcal{NP}$ closely connects with the field of proof complexity. In particular, we have shown that no Frege proof systems are polynomial bounded (Theorem 7), thus answering an open problem in [Pud08]. As pointed out in Remark 8.2, we do the opposite direction of Cook’s program. That is, we first prove that $\mathcal{NP}\neq{\rm co}\mathcal{NP}$, and then apply it to prove the lower bounds on the length of proofs in Frege proof systems. Note that one of the key points to show the Theorem 7 is that nondeterministic Turing machines can simulate Frege proofs. But, we do not know whether the nondeterministic Turing machines can simulate Extended Frege proofs. For more details about Extended Frege Systems, please consult the reference [CR79, CN10, Kra95].

### v8

@@ Abstract
We prove in this paper that there is a language $L_{d}$ accepted by some nondeterministic Turing machines but not by any ${\rm co}\mathcal{NP}$-machines (defined later). Then we further show that $L_{d}$ is in $\mathcal{NP}$, thus proving that $\mathcal{NP}\neq{\rm co}\mathcal{NP}$. The main techniques used in this paper are lazy-diagonalization and the novel new technique developed in the author’s recent work [Lin21]. Further, since there exists some oracle $A$ such that $\mathcal{NP}^{A}={\rm co}\mathcal{NP}^{A}$, we then explore what mystery behind it and show that if $\mathcal{NP}^{A}={\rm co}\mathcal{NP}^{A}$ and under some rational assumptions, the set of all polynomial-time co-nondeterministic oracle Turing machines with oracle $A$ is not enumerable, thus showing that the technique of lazy-diagonalization is not applicable for the first half of the whole step to separate $\mathcal{NP}^{A}$ from ${\rm co}\mathcal{NP}^{A}$. As a by-product, we reach the important result that $\mathcal{P}\neq\mathcal{NP}$ [Lin21] once again, which is clear from the above outcome and the well-known fact that $\mathcal{P}={\rm co}\mathcal{P}$. Next, we show that the complexity class ${\rm co}\mathcal{NP}$ has intermediate languages, i.e., there exists language $L_{inter}\in{\rm co}\mathcal{NP}$ which is not in $\mathcal{P}$ and not ${\rm co}\mathcal{NP}$-complete. We also summarize other direct consequences implied by our main outcome such as $\mathcal{NEXP}\neq{\rm co}\mathcal{NEXP}$ and other which belongs to the area of proof complexity. Lastly, we show a lower bounds result for Frege proof systems, i.e., no Frege proof systems can be polynomially bounded.
@@ 8.1 Lower Bounds for Frege Systems
The following theorem gives a necessary and sufficient condition for the complexity class $\mathcal{NP}$ being closed under complementation. For completeness, the proof is also quoted from [CR79]:
@@ 9 Conclusions
thus resolving a very important and long-standing conjecture in computational complexity theory. Specifically, we first showed that the set of all ${\rm co}\mathcal{NP}$-machines is enumerable. Then, we constructed a universal nondeterministic Turing machine $D$ to simulate all ${\rm co}\mathcal{NP}$-machines and to do the lazy-diagonalization operation to create a language $L_{d}$ which is not in ${\rm co}\mathcal{NP}$. Next, we applied the technique developed in the author’s recent work [Lin21] to show that the language $L_{d}$ accepted by the universal nondeterministic Turing machine $D$ is in fact in $\mathcal{NP}$, thus establishing the desired main conclusion of this paper.
We stress again that one of the important techniques used to prove lower bounds (i.e., Theorem 4.1) is the so-called lazy-diagonalization [FS07, AB09], and the essential technique used to show the upper bounds result, i.e., the result that the language $L_{d}\in\mathcal{NP}$, is essentially the same as the one developed in the author’s recent work [Lin21].
We have broken the so-called Relativization Barrier. Since by the result of [BGS75], it is not hard to show that there exists some oracle $A$ such that
@@
which suggests that ordinary diagonalization techniques (lazy-diagonalization) are not capable of proving $\mathcal{NP}\neq{\rm co}\mathcal{NP}$, see e.g. [BGS75]. We explored the mystery behind this phenomenon by proving that if $\mathcal{NP}^{A}={\rm co}\mathcal{NP}^{A}$ and under some rational assumptions, then the set of all ${\rm co}\mathcal{NP}$-machines with oracle $A$ is in fact not enumerable, thus showing that the ordinary diagonalization techniques (i.e., lazy-diagonalization) will generally not apply to the relativized versions of the $\mathcal{NP}$ versus ${\rm co}\mathcal{NP}$ problem.
As a by-product, the important result that $\mathcal{P}\neq\mathcal{NP}$ which is shown in [Lin21], also follows. Furthermore, the conclusion of $\mathcal{NEXP}\neq{\rm co}\mathcal{NEXP}$ follows from the similar proof of that $\mathcal{NP}\neq{\rm co}\mathcal{NP}$. The above two outcomes can be seen as, in a sense, the direct or indirect consequences of the proof procedures of
$\mathcal{NP}\neq{\rm co}\mathcal{NP}.$
Note again that by Corollary 3 we in fact have shown $\mathcal{P}\neq{\rm co}\mathcal{NP}$ because if $\mathcal{P}={\rm co}\mathcal{NP}$, then $\mathcal{P}={\rm co}\mathcal{P}=\mathcal{NP}$, a contradiction to Corollary 3. As a matter of fact and interestingly, we have also shown by Theorem 5 that there exists a ${\rm co}\mathcal{NP}$-intermediate language in ${\rm co}\mathcal{NP}$ which is not in $\mathcal{P}$ nor ${\rm co}\mathcal{NP}$-complete, i.e., the complexity class ${\rm co}\mathcal{NP}$ also has a rich structure. We have also obtained other interesting consequences, such as Corollary 6 which states that there is no super proof system.
As mentioned in the introduction Section, the problem of $\mathcal{NP}$ versus ${\rm co}\mathcal{NP}$ closely connects with the field of proof complexity. In particular, we have shown that no Frege proof systems are polynomially bounded (Theorem 7), thus answering an open problem in [Pud08]. As pointed out in Remark 8.2, we do the opposite direction of Cook’s program. That is, we first prove that $\mathcal{NP}\neq{\rm co}\mathcal{NP}$, and then apply it to prove the lower bounds on the length of proofs in Frege proof systems. Note that one of the key points to show the Theorem 7 is that nondeterministic Turing machines can simulate Frege proofs (see [Bus99]). But we do not know whether the nondeterministic Turing machines can simulate Extended Frege proofs. For more details about Extended Frege Systems, please consult the reference [CR79, CN10, Kra95].